How much salt must be evaporated for 120 kg of solution which is 3% salt to make a solution of 5% salt?
Answer by stanbon(75887)

How much water must be evaporated for 120 kg of solution which is 3% salt to make a solution of 5% salt?
======
Equation:l
water - water = water
0.97*120 - 1x = 0.95(120-x)
97(120)- 100X = 95*120-95X
2(120)=5X
X = 48 kg (amt. of water that must be evaporated)
Cheers,
Stan H.